OpenStudy (anonymous):

how would I solve 10095 m/s to miles/s

OpenStudy (anonymous):

Do you have a good conversion between metric and "regular" (like feet or miles)?

OpenStudy (anonymous):

Or meters to yards?

OpenStudy (anonymous):

i know theres 5280 feet in a mile and 1760 meters in a mile

OpenStudy (anonymous):

yep, right idea. Do you have anything that says something like centimeters per inch or yards per meter?

OpenStudy (anonymous):

yes 2.54 centimeters= 1 inch

OpenStudy (anonymous):

ok, well, that will work. It's just a bunch of step-by-step unit conversions...

OpenStudy (anonymous):

\[10095 * \frac{ m }{ s } * 100 \frac{ cm }{ m } * \frac{ 1 inch }{ 2.54 cm } * \frac{ 1 foot }{12 inches} * \frac{ 1 mile}{ 5280 feet } = 6.2727 miles/second \]

OpenStudy (anonymous):

you just have to keep straight cancelling out the unit you don't want in each step by putting a conversion factor, in the form of a fraction with the right units on top and bottom. Keep converting till you get to the real units you want at the end.

OpenStudy (anonymous):

Sometimes it helps to write out what you're given (in this case, m/s) on the left, and what you need (miles/s) on the right, and then start filling in conversion factors one at a time to "connect the dots" from m/s to miles/s

OpenStudy (anonymous):

oooooh ok and the seconds i dont do anything with since there both divided seconds?

OpenStudy (anonymous):

right on! But lots of times you want speed in terms of miles per hour, right? So, watch out on a test question! You might ALSO have to put in even more conversions to convert seconds to minutes and minutes to hours

OpenStudy (anonymous):

ok got it! preciate it

OpenStudy (anonymous):

No problem, glad to help :)